Lean Accounting

Vintage Audio Inc. manufactures audio speakers. Each speaker requires $119 per unit of direct materials. The speaker manufacturing assembly cell includes the following estimated costs for the period:

Speaker assembly cell, estimated costs:
Labor $80,070
Depreciation 10,740
Supplies 3,910
Power 2,930
Total cell costs for the period $97,650

The operating plan calls for 175 operating hours for the period. Each speaker requires 10 minutes of cell process time. The unit selling price for each speaker is $322. During the period, the following transactions occurred:

  1. Purchased materials to produce 435 speaker units.
  2. Applied conversion costs to production of 415 speaker units.
  3. Completed and transferred 395 speaker units to finished goods.
  4. Sold 380 speaker units.

There were no inventories at the beginning of the period.

a. Journalize the summary transactions (1)-(4) for the period. Do not round interim calculations.

1.
2.
3.
4. Sale
4. Cost

b. Determine the ending balance of raw and in process inventory and finished goods inventory.

Raw and In Process Inventory, ending balance $
Finished Goods Inventory, ending balance $
Answer #1
Budgeted cell conversion cost per hour 558 per hour =97650/175
Budgeted cell conversion cost per unit 93.00 per unit =558*(10/60)
a
Raw and in process inventory 51765 =435*119
     Accounts Payable 51765
Raw and in process inventory 38595 =415*93
       Conversion costs 38595
Finished goods inventory 83740 =395*(119+93)
       Raw and in process inventory 83740
Accounts Receivable 122360 =380*322
       Sales 122360
Cost of goods sold 80560 =380*(119+93)
        Finished goods inventory 80560
b
Raw and In Process Inventory, ending balance 6620 =51765+38595-83740
Finished Goods Inventory, ending balance 3180 =83740-80560